# Network inference via approximate Bayesian computation. Illustration on a stochastic multi-population neural mass model

### v3

@@ 3.3.2 Algorithm implementation details
The number $M$ of kept particles per iteration is set to $500$. The initial threshold $\delta_{1}$ for the first iteration is obtained by a reference table acceptance-rejection ABC pilot run. Under the given prior, we produce $10^{4}$ distances and then choose $\delta_{1}$ as their median. For $r>1$, the threshold $\delta_{r}$ is chosen as the median of the $M$ distances computed at the previous iteration if the acceptance rate of particles at the previous iteration is larger than 1%. Otherwise $\delta_{r}$ is chosen as the $75$-th percentile of the $M$ distances computed at iteration $r-1$ (to enable further iterations of the algorithm at its final stage). The algorithm is stopped after the acceptance rate has dropped below the prefixed threshold of 0.1%.
@@ 4.1 Parameter vector and prior distribution
We aim to infer both the continuous and the binary parameters of the stochastic multi-population JR-NMM. The continuous parameters consist of the intrinsic model parameters $A_{k}>0$, $k=1,\ldots,N$, which play a central role in the (non-)activation of neural populations (cf. Section 2), as well as the coupling strength parameters $L>0$ and $10\ll c<$ 1, see (5) and (6). The constants $\mu_{k}$ and $\sigma_{k}$ are fixed to $90$ and $500$, respectively, and the remaining continuous constants are fixed according to the standard values reported in Table 1. The binary parameters consist of the coupling direction parameters $\rho_{jk}$, $j,k=1,\ldots,N$, $j\neq k$, see (5). Thus, our goal is to apply the nSMC-ABC Algorithm 1 for inference of the $(N+2+N(N-1))$-dimensional parameter vector
@@ 4.2 Description of observed data
$\begin{split}\theta&=(A_{1},A_{2},A_{3},A_{4},L,\textrm{vec}(\mathcal{P}))=(3.6,3.25,3.25,3.25,700,\textrm{vec}(\mathcal{P})),\\ &\ \text{with }\rho_{12}=\rho_{23}=\rho_{34}=1\text{ and all other }\rho_{jk}=0\mbox{ (cascade)},\\ \theta&=(A_{1},A_{2},A_{3},A_{4},L,c,\textrm{vec}(\mathcal{P}))=(3.6,3.25,3.25,3.25,700,0.8,\textrm{vec}(\mathcal{P})),\\ &\ \text{with }\rho_{12}=\rho_{23}=\rho_{34}=\rho_{13}=\rho_{32}=1\text{ and all other }\rho_{jk}=0\mbox{ (partially connected)},\\ \theta&=(A_{1},A_{2},A_{3},A_{4},L,c,\textrm{vec}(\mathcal{P}))=(3.25,3.25,3.25,3.25,700,0.8,\textrm{vec}(\mathcal{P})),\\ &\ \text{with }\rho_{jk}=1\text{ for all }j,k\mbox{ (fully connected)}.\end{split}$ (17)
@@ 4.3 Estimation results
In Table 2, we report the estimates $\hat{\rho}_{jk}$ of $\rho_{jk}$ obtained as ABC posterior modes, together with the ABC posterior means in parentheses. All estimates coincide with the underlying true values of the $\rho_{jk}$ parameters in (17), the full network thus being correctly identified for all scenarios. In most cases (except for $\rho_{13}$ and $\rho_{14}$ in the partially connected scenario), even the posterior means coincide with the true values for $\rho_{jk}$ being equal to 00 or $1$, respectively.
@@ 5.2 Inference
Denote the channels FP1-F7, FP1-F3, FP2-F4 and FP2-F8 by Population $1$, $2$, $3$ and $4$, respectively. As the distance between channels on different hemispheres is larger than that on the same hemisphere, the distance between the channels FP1-F3 and FP2-F4 is larger than that between FP1-F7 and FP1-F3 or between FP2-F4 and FP2-F8. Similar to (6), we therefore assume the following matrix of coupling strength parameters
@@
with unknown parameters $L>0$ and $10\ll c<$ 1.
@@
We assume the continuous model parameters, i.e., the activation parameters $A_{k}$, $k=1,\ldots,4$, the noise intensity parameters $\sigma_{L}:=\sigma_{1}=\sigma_{2}$ (left hemisphere) and $\sigma_{R}:=\sigma_{3}=\sigma_{4}$ (right hemisphere) as well as the input parameters $\mu_{L}:=\mu_{1}=\mu_{2}$ (left hemisphere) and $\mu_{R}:=\mu_{3}=\mu_{4}$ (right hemisphere) to be unknown. For the remaining parameters, we use the standard values reported in Table 1, except for setting $b=20$ and $C=70$ (values chosen based on pilot experiments). We thus aim to infer the $(10+12)$-dimensional parameter vector
@@
In Table 3, we report estimates (values in $\{0,1\}$ obtained as marginal posterior modes) for the $12$ network parameters $\rho_{jk}$, $j,k=1,\ldots,4$, $j\neq k$, as well as the corresponding posterior means in parentheses. All posterior means for the during seizure scenarios and most posterior means for the before seizure scenarios lie outside the interval $[1/3,2/3]$, being thus “close” to 00 or $1$, respectively. Therefore, the results suggest clear connectivity structures, with the estimated networks visualized in Figure 7. Remarkably, we observe similarities in the inferred networks for both datasets. For example, the estimation results suggest a stronger connectivity in the left brain hemisphere during seizure, compared to before seizure (in particular, for both datasets the connection from Population 2 to Population 1 is only present during seizure). Moreover, for both datasets we observe less connectivity in the right brain hemisphere as well as less connectivity from the right to the left hemisphere during seizure, compared to before seizure. There are also differences among the two datasets though, e.g., the connection from Population 3 to Population 4 is only present for the second dataset.
@@
Figure 7: Inferred networks from EEG recordings. Estimated networks from two recordings with $N=4$ channels of an 11 year old female patient before and during epileptic seizure. Populations 1, 2, 3, and 4 refer to channels FP1-F7, FP1-F3, FP2-F4, and FP2-F8, respectively. The dotted orange connections are estimated with a posterior mean within $[1/3,2/3]$ (cf. Table 3).
@@ 6 Conclusion and discussion
To address these challenges, we employ the likelihood-free ABC methodology, introducing the nSMC-ABC method, an adaptation of the gold-standard SMC-ABC algorithm. Our method leverages the introduced $\{0,1\}$-valued network indicators, allowing for a reduction of computationally expensive continuous parameters and enabling the use of an efficient perturbation kernel that independently updates continuous and binary parameters. Two key algorithm components are essential: a reliable and computationally efficient numerical simulation method for synthetic data generation, and informative summary statistics capturing model dynamics independent of the underlying intrinsic stochasticity. To this end, we further develop the ABC approach of [7] to multiple interacting populations. First, we construct a numerical scheme based on the splitting approach. It exploits the model’s Hamiltonian structure and scales automatically with $N$. Standard methods like Euler-Maruyama would either fail in estimation or make the algorithm computationally infeasible due to the requirement of very small time discretization steps (see [7] for an illustration of this issue on the stochastic JR-NMM with $N=1$). Second, we propose summaries that incorporate both individual signal features and pairwise connectivity measures among signals by mapping the $N$-dimensional time series to the $N$ corresponding densities and spectral densities, and the $N(N-1)$ cross-correlation functions.
@@ A.2 Construction of the numerical splitting method for the stochastic multi-population JR-NMM
[…]
@@ Appendix B Deeper investigation of the proposed nSMC-ABC algorithm
Imagine that the binary coupling direction parameters $\rho_{jk}\in\{0,1\}$ were absent in (5). Then, the continuous coupling strength parameters $K_{jk}\geq 0$ would have to be used to determine the network structure and a standard SMC-ABC algorithm would be required for their inference. This would be computationally expensive, since $N(N-1)$ continuous network parameters would have to be inferred. Instead, the use of binary network indicators allows to reduce the number of continuous network parameters (here, to only two parameters $L$ and $c$, see (6)), which significantly reduces the number of model simulations required in nSMC-ABC to reach the desired posterior region, and thus the computational cost of the algorithm. In addition, for the standard SMC-ABC algorithm a criterion would be required to decide whether there is or not a connection from population $j$ to $k$ (e.g., the marginal posterior median of $K_{jk}$ being larger than some threshold value). By using binary indicators and nSMC-ABC, we avoid setting such decision criterion, which brings an extra level of arbitrariness. Therefore, using binary indicators and nSMC-ABC has two main advantages. First, it allows to reduce the number of continuous parameters, and thus the computational cost of the algorithm. Second, it does not require a decision criterion for the presence or absence of network connections.
@@
where $\hat{v}_{u,r}$ is the population variance obtained from the previous population of that binary parameter at iteration $r-1$. The larger the variance in the previous population of a binary parameter, the larger the probability that the sampled value is flipped (i.e., the less we trust the sampled value). This is in agreement with the perturbation of the continuous parameters, in the sense that the larger the variance in the previous population, the “stronger” the perturbation is.
@@
We use the population variance in (24), because, in contrast to the sample variance, it is always between 00 and $1/4$, guaranteeing $q_{\textrm{flip}}$ to be a probability value within $[0,1]$.
We apply the nSMC-ABC method based on this alternative perturbation kernel to the same observed data as before, and report the results as orange dotted-dashed lines in Figures 13-15. We observe that also this version of the algorithm leads to similar marginal posteriors as nSMC-ABC using the proposed perturbation kernel, requiring significantly more model simulations to reach the desired parameter regions though. This becomes evident when looking at Figure 13 and 14, where we report the $90$% CIs of the marginal posterior densities of the continuous parameters and the marginal posterior means of the $\rho_{jk}$-parameters, respectively, as functions of the number of model simulations. Moreover, the distance-threshold $\delta$ decreases slower in the alternative nSMC-ABC algorithm (cf. left panel of Figure 15), which requires more model simulations (and thus more computational time) to correctly infer the network (cf. the F1-scores in the middle and right panel of Figure 15). This indicates that the proposed nSMC-ABC method yields satisfactory inferential results for different types of perturbation kernels, while benefiting from an independent treatment of the continuous and binary parameters, as proposed in Algorithm 1.
@@ C.1 Impact of the choice of the coupling strength structure
These findings are confirmed by looking at Figure 17, where we report the corresponding marginal posterior means of the binary coupling direction parameters $\rho_{jk}$, $j,k=1,\ldots,4$, $j\neq k$, as functions of the number of model simulations, for the before (dashed lines) and during (solid lines) seizure periods of data1 (cf. Figure 5a). The respective blue, red and black lines correspond to $K$ (18), $\bar{K}$ and $\widetilde{K}$ (25), respectively. Also for the binary parameters, the inferential results are similar under the three different coupling structures, leading to the same inferred networks. Moreover, for some of the network parameters, we observe a faster convergence of the algorithm (to the true parameter values) in the during seizure setting. This is particularly evident for $\rho_{21}$, but also, for example, for $\rho_{31}$, $\rho_{34}$ and $\rho_{43}$. Note also that the results for $\rho_{42}$ and $\rho_{43}$ under the before seizure scenario confirm the low evidence that we have for those connections to be present or not, as indicated with the dotted orange lines in Figure 7a (left network).
@@ C.2 Inference from another patient’s EEG recordings
In this section, we investigate two EEG recordings of the second pediatric subject from the CHB-MIT Scalp EEG Database, an $11$ year old male patient (see Section 5.1 of the main manuscript for the data details). The two investigated datasets are recorded over time intervals of length $164$ and $162$ seconds, respectively. They are available in the edf-files chb02_16 and chb02_16+, and denoted by data1 and data2 in the following. A visualization of these datasets is provided in Figure 19, where the dotted red lines separate the data into the respective period before and during seizure (see the seizure classification in [35]). As in Section 5.2 of the main manuscript, we denote the channels FP1-F7, FP1-F3, FP2-F4, and FP2-F8 by Population $1$, $2$, $3$, and $4$, respectively, and apply the proposed nSMC-ABC Algorithm 1 (under the coupling strength structure $K$ (18)) to fit the stochastic multi-population JR-NMM to these four EEG segments.
@@
Figure 21: Inferred networks from EEG recordings. Estimated networks from two recordings with $N=4$ channels of an 11 year old male patient before and during epileptic seizure. Populations 1, 2, 3, and 4 refer to channels FP1-F7, FP1-F3, FP2-F4, and FP2-F8, respectively. The dotted orange connections are estimated with a posterior mean within $[1/3,2/3]$ (cf. Table 4).
In Table 4, we report the network estimates of the binary $\rho_{jk}$-parameters and the corresponding marginal posterior means in parentheses. Similarly to the first subject, all posterior means for the during seizure scenario are either close to $1$ or 00, clearly indicating whether a connection is present or not. For the before seizure scenarios, only the posterior means of $\rho_{14}$ for data1 and $\rho_{31}$ for data2 lie inside $[1/3,2/3]$, the possible unclear connections marked as dotted orange lines in Figure 21, where we visualize the inferred networks. As for the first patient, we observe some similarities across data1 and data2. For example, in both recordings, we observe a very clear activation of $\rho_{12}$ in the left brain hemisphere and of $\rho_{43}$ in the right brain hemisphere during seizure, connections which were both not present before seizure. Moreover, there is less connectivity between the two hemispheres during seizure, both from left to right and from right to left.

### v4

@@ 3.3.2 Algorithm implementation details
The number $M$ of kept particles per iteration is set to $500$. The initial threshold $\delta_{1}$ for the first iteration is obtained by a reference table acceptance-rejection ABC pilot run. Under the given prior, we produce $10^{4}$ distances and then choose $\delta_{1}$ as their median. For $r>1$, the threshold $\delta_{r}$ is chosen as the median of the $M$ distances computed at the previous iteration if the acceptance rate of particles at the previous iteration is larger than 1%. Otherwise $\delta_{r}$ is chosen as the $75$-th percentile of the $M$ distances computed at iteration r1r-1 (to enable further iterations of the algorithm at its final stage). The algorithm is stopped after the acceptance rate has dropped below the prefixed threshold of 0.1%.
@@ 4.1 Parameter vector and prior distribution
We aim to infer both the continuous and the binary parameters of the stochastic multi-population JR-NMM. The continuous parameters consist of the intrinsic model parameters $A_{k}>0$, $k=1,\ldots,N$, which play a central role in the (non-)activation of neural populations (cf. Section 2), as well as the coupling strength parameters $L>0$ and $0\ll c<1$, see (5) and (6). The constants $\mu_{k}$ and $\sigma_{k}$ are fixed to $90$ and $500$, respectively, and the remaining continuous constants are fixed according to the typical values reported in Table 1. The binary parameters consist of the coupling direction parameters $\rho_{jk}$, $1,\ldots,N, jkj\neq k$, see (5). Thus, our goal is to apply the nSMC-ABC Algorithm 1 for inference of the (N+2+N(N1))(N+2+N(N-1))-dimensional parameter vector
@@ 4.2 Description of observed data
[…]
@@ 4.3 Estimation results
In Table 2, we report the estimates $\hat{\rho}_{jk}$ of $\rho_{jk}$ obtained as ABC posterior modes, together with the ABC posterior means in parentheses. All estimates coincide with the underlying true values of the $\rho_{jk}$ parameters in (17), the full network thus being correctly identified for all scenarios. In most cases (except for $\rho_{13}$ and $\rho_{14}$ in the partially connected scenario), even the posterior means coincide with the true values for $\rho_{jk}$ being equal to 0 or $1$, respectively.
@@ 5.2 Inference
Denote the channels FP1-F7, FP1-F3, FP2-F4 and FP2-F8 by Population $1$, 22, 33 and $4$, respectively. As the distance between channels on different hemispheres is larger than that on the same hemisphere, the distance between the channels FP1-F3 and FP2-F4 is larger than that between FP1-F7 and FP1-F3 or between FP2-F4 and FP2-F8. Similar to (6), we therefore assume the following matrix of coupling strength parameters
@@
with unknown parameters $L>0$ and $0\ll c<1$.
@@
We assume the continuous model parameters, i.e., the activation parameters $A_{k}$, $k=1,\ldots,4$, the noise intensity parameters $\sigma_{L}:=\sigma_{1}=\sigma_{2}$ (left hemisphere) and $\sigma_{R}:=\sigma_{3}=\sigma_{4}$ (right hemisphere) as well as the input parameters $\mu_{L}:=\mu_{1}=\mu_{2}$ (left hemisphere) and $\mu_{R}:=\mu_{3}=\mu_{4}$ (right hemisphere) to be unknown. For the remaining parameters, we use the typical values reported in Table 1, except for setting $b=20$ and $C=70$ (values chosen based on pilot experiments). We thus aim to infer the $(10+12)$-dimensional parameter vector
@@
In Table 3, we report estimates (values in $\{0,1\}$ obtained as marginal posterior modes) for the $12$ network parameters $\rho_{jk}$, $1,\ldots,4, jkj\neq k$, as well as the corresponding posterior means in parentheses. All posterior means for the during seizure scenarios and most posterior means for the before seizure scenarios lie outside the interval $[1/3,2/3]$, being thus “close” to 0 or $1$, respectively. Therefore, the results suggest clear connectivity structures, with the estimated networks visualized in Figure 7. Remarkably, we observe similarities in the inferred networks for both datasets. For example, the estimation results suggest a stronger connectivity in the left brain hemisphere during seizure, compared to before seizure (in particular, for both datasets the connection from Population 2 to Population 1 is only present during seizure). Moreover, for both datasets we observe less connectivity in the right brain hemisphere as well as less connectivity from the right to the left hemisphere during seizure, compared to before seizure. There are also differences among the two datasets though, e.g., the connection from Population 3 to Population 4 is only present for the second dataset.
@@
Figure 7: Inferred networks from EEG recordings. Estimated networks from two recordings with $N=4$ channels of an $1$ year old female patient before and during epileptic seizure. Populations 1, 2, 3, and 4 refer to channels FP1-F7, FP1-F3, FP2-F4, and FP2-F8, respectively. The dotted orange connections are estimated with a posterior mean within $[1/3,2/3]$ (cf. Table 3).
@@ 6 Conclusion and discussion
To address these challenges, we employ the likelihood-free ABC methodology, introducing the nSMC-ABC method, an adaptation of the gold-standard SMC-ABC algorithm. Our method leverages the introduced $\{0,1\}$-valued network indicators, allowing for a reduction of computationally expensive continuous parameters and enabling the use of an efficient perturbation kernel that independently updates continuous and binary parameters. Two key algorithm components are essential: a reliable and computationally efficient numerical simulation method for synthetic data generation, and informative summary statistics capturing model dynamics independent of the underlying intrinsic stochasticity. To this end, we further develop the ABC approach of [8] to multiple interacting populations. First, we construct a numerical scheme based on the splitting approach. It exploits the model’s Hamiltonian structure and scales automatically with $N$. Standard methods like Euler-Maruyama would either fail in estimation or make the algorithm computationally infeasible due to the requirement of very small time discretization steps (see [8] for an illustration of this issue on the stochastic JR-NMM with $N=1$). Second, we propose summaries that incorporate both individual signal features and pairwise connectivity measures among signals by mapping the $N$-dimensional time series to the $N$ corresponding densities and spectral densities, and the N(N1)N(N-1) cross-correlation functions.
@@ A.2 Construction of the numerical splitting method for the stochastic multi-population JR-NMM
$\textrm{Cov}(\Delta)=\int\limits_{0}^{\Delta}e^{F(\Delta-s)}\Sigma_{0}\Sigma_{0}^{\top}\left(e^{F(\Delta-s)}\right)^{\top}ds.$
@@ Appendix B Deeper investigation of the proposed nSMC-ABC algorithm
Imagine that the binary coupling direction parameters $\rho_{jk}\in\{0,1\}$ were absent in (5). Then, the continuous coupling strength parameters $K_{jk}\geq 0$ would have to be used to determine the network structure and a standard SMC-ABC algorithm would be required for their inference. This would be computationally expensive, since N(N1)N(N-1) continuous network parameters would have to be inferred. Instead, the use of binary network indicators allows to reduce the number of continuous network parameters (here, to only two parameters $L$ and $c$, see (6)), which significantly reduces the number of model simulations required in nSMC-ABC to reach the desired posterior region, and thus the computational cost of the algorithm. In addition, for the standard SMC-ABC algorithm a criterion would be required to decide whether there is or not a connection from population $j$ to $k$ (e.g., the marginal posterior median of $K_{jk}$ being larger than some threshold value). By using binary indicators and nSMC-ABC, we avoid setting such decision criterion, which brings an extra level of arbitrariness. Therefore, using binary indicators and nSMC-ABC has two main advantages. First, it allows to reduce the number of continuous parameters, and thus the computational cost of the algorithm. Second, it does not require a decision criterion for the presence or absence of network connections.
@@
where $\hat{v}_{u,r}$ is the population variance obtained from the previous population of that binary parameter at iteration r1r-1. The larger the variance in the previous population of a binary parameter, the larger the probability that the sampled value is flipped (i.e., the less we trust the sampled value). This is in agreement with the perturbation of the continuous parameters, in the sense that the larger the variance in the previous population, the “stronger” the perturbation is.
@@
We use the population variance in (24), because, in contrast to the sample variance, it is always between 0 and $1/4$, guaranteeing $q_{\textrm{flip}}$ to be a probability value within $[0,1]$.
We apply the nSMC-ABC method based on this alternative perturbation kernel to the same observed data as before, and report the results as orange dotted-dashed lines in Figures 13-15. We observe that also this version of the algorithm leads to similar marginal posteriors as nSMC-ABC using the proposed perturbation kernel, requiring significantly more model simulations to reach the desired parameter regions though. This becomes evident when looking at Figure 13 and 14, where we report the 9090% CIs of the marginal posterior densities of the continuous parameters and the marginal posterior means of the $\rho_{jk}$-parameters, respectively, as functions of the number of model simulations. Moreover, the distance-threshold $\delta$ decreases slower in the alternative nSMC-ABC algorithm (cf. left panel of Figure 15), which requires more model simulations (and thus more computational time) to correctly infer the network (cf. the F1-scores in the middle and right panel of Figure 15). This indicates that the proposed nSMC-ABC method yields satisfactory inferential results for different types of perturbation kernels, while benefiting from an independent treatment of the continuous and binary parameters, as proposed in Algorithm 1.
@@ C.1 Impact of the choice of the coupling strength structure
These findings are confirmed by looking at Figure 17, where we report the corresponding marginal posterior means of the binary coupling direction parameters $\rho_{jk}$, $1,\ldots,4, jkj\neq k$, as functions of the number of model simulations, for the before (dashed lines) and during (solid lines) seizure periods of data1 (cf. Figure 5a). The respective blue, red and black lines correspond to $K$ (18), $\bar{K}$ and $\widetilde{K}$ (25), respectively. Also for the binary parameters, the inferential results are similar under the three different coupling structures, leading to the same inferred networks. Moreover, for some of the network parameters, we observe a faster convergence of the algorithm under $K$ (18) (to 0/1) in the during seizure setting. This is particularly evident for $\rho_{21}$, but also, for example, for $\rho_{31}$, $\rho_{34}$ and $\rho_{43}$. Note also that the results for $\rho_{42}$ and $\rho_{43}$ under the before seizure scenario confirm the low evidence that we have for those connections to be present or not, as indicated with the dotted orange lines in Figure 7a (left network).
@@ C.2 Inference from another patient’s EEG recordings
In this section, we investigate two EEG recordings of the second pediatric subject from the CHB-MIT Scalp EEG Database, an $11$ year old male patient (see Section 5.1 of the main manuscript for the data details). The two investigated datasets are recorded over time intervals of length $164$ and $162$ seconds, respectively. They are available in the edf-files chb02_16 and chb02_16+, and denoted by data1 and data2 in the following. A visualization of these datasets is provided in Figure 19, where the dotted red lines separate the data into the respective period before and during seizure (see the seizure classification in [36]). As in Section 5.2 of the main manuscript, we denote the channels FP1-F7, FP1-F3, FP2-F4, and FP2-F8 by Population $1$, 22, 33, and $4$, respectively, and apply the proposed nSMC-ABC Algorithm 1 (under the coupling strength structure $K$ (18)) to fit the stochastic multi-population JR-NMM to these four EEG segments.
@@
Figure 21: Inferred networks from EEG recordings. Estimated networks from two recordings with $N=4$ channels of an $1$ year old male patient before and during epileptic seizure. Populations 1, 2, 3, and 4 refer to channels FP1-F7, FP1-F3, FP2-F4, and FP2-F8, respectively. The dotted orange connections are estimated with a posterior mean within $[1/3,2/3]$ (cf. Table 4).
In Table 4, we report the network estimates of the binary $\rho_{jk}$-parameters and the corresponding marginal posterior means in parentheses. Similarly to the first subject, all posterior means for the during seizure scenario are either close to $1$ or 0, clearly indicating whether a connection is present or not. For the before seizure scenarios, only the posterior means of $\rho_{14}$ for data1 and $\rho_{31}$ for data2 lie inside $[1/3,2/3]$, the possible unclear connections marked as dotted orange lines in Figure 21, where we visualize the inferred networks. As for the first patient, we observe some similarities across data1 and data2. For example, in both recordings, we observe a very clear activation of $\rho_{12}$ in the left brain hemisphere and of $\rho_{43}$ in the right brain hemisphere during seizure, connections which were both not present before seizure. Moreover, there is less connectivity between the two hemispheres during seizure, both from left to right and from right to left.